assume that tom and jerry from a partnership with investment of 88000$ and120000$ separately. Determine the partnership in net income of 88000$ for the division of the partnership net income under the following conditions.

  1. No agreement concerning division
  2. Division is made in the ration of their original investment
  3. Division is made according to the agreement that he rate of 8% allowed on original investment and the remainder divided in the ratio of 2:1
1
Expert's answer
2021-12-05T18:56:47-0500

(1)

No agreement concerning division:

Division of partnership will be equal, thus,

"=0.5\times88000=44000"

Tom and Jerry will each get $4400.

(2)

According to ration of original investment:

Total Investment,

"=88000+120000=208000"

Tom will get:

"\frac{88000}{208000}\times88000=37230.77"

Jerry will get:

"\frac{120000}{208000}\times88000=50769.23"

(3)

Rate of 8% on original investment:

Tom,"=\frac{8}{100}\times88000=7040"

Jerry,"=\frac{8}{100}\times120000=9600"

Remainder divided in the ration 2:1

Remainder"=88000-(7040+9600)"

"=71360"

For Tom:

"=\frac{2}{3}\times71360=47573.33"

For Jerry:

"=\frac{1}{3}\times71360=23786.67"

Total share:

Tom:"=7040+47573.33=54613.33"

=$54613.33

Jerry:

"=9600+23786.67=33386.67"

=$33386.67